OpenStudy (anonymous):

Solve the following system of equations for x: 2x+y=4 x-2y=7 1)2 2)3 3)5 4)7

OpenStudy (saifoo.khan):

3

OpenStudy (anonymous):

you can use substitution to solve first for y then for x. Just get one equation to look like x=.... then solve for x with substitution

OpenStudy (anonymous):

man this is easy multiply the 2nd eq with 2 you will get 2x-4y=14 then minus them 2x+y=4 - 2x-4y=14 5y = 10 y = 2 x = 7-(-(2))

OpenStudy (anonymous):

= 5 ..3

OpenStudy (anonymous):

it's 3 sorry I'm wrong. 2x(2x+y=4) 4x+2y =8 x-2y = 7 add them 5x = 15 x = 3